### v1

# EFRame: Deeper Reasoning via Exploration-Filtering-Replay Reinforcement Learning Framework

Abstract
Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs). Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks. To address these limitations, we introduce EFRame, an Exploration-Filtering-Replay framework that systematically augments GRPO along three critical dimensions. EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples. EFRame establishes a complete and stable learning cycle, guiding the model through a structured transition from exploration to convergence. Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO. Furthermore, EFRame enables a more fine-grained categorization of training samples, allowing for a deeper analysis of how different types of samples contribute to the learning process in RL. Our code is avail able at https://github.com/597358816/EFRame.
1 Introduction
Reinforcement learning (RL) has played a pivotal role in the post-training of large language models (LLMs) GLM et al. (2024); Touvron et al. (2023). In the early stages, approaches such as Proximal Policy Optimization (PPO) and Direct Preference Optimization (DPO) were widely adopted under the Reinforcement Learning from Human Feedback (RLHF), effectively aligning model outputs with human preferences through preference-based rewards Schulman et al. (2017); Rafailov et al. (2023); Zhong et al. (2024); Wang et al. (2024b). More recently, the focus has shifted towards Reinforcement Learning with Verifiable Rewards (RLVR), which seeks to provide more consistent and scalable supervision by leveraging structured, verifiable reward signals Mroueh (2025). Within this paradigm, DeepSeek-R1 introduced Group Relative Policy Optimization (GRPO), a lightweight and efficient variant of PPO that compute relative rewards within rollout groups to eliminate the need for the critic model Shao et al. (2024); Liu et al. (2024); Guo et al. (2025). GRPO has demonstrated remarkable improvements in training efficiency and has shown promising potential in enhancing the reasoning capabilities of LLMs.
RLVR relies on extracting reward signals from the model’s own pass@$k$ outputs to improve its pass@1 performance. However, this design inherently couples final performance to the choice of $k$, as the reward signal vanishes when all sampled responses are either correct or incorrect, making it impossible for the model to extract useful knowledge from such data. This constrains RLVR’s learning capacity to the base model’s pass@$k$, limiting further exploration and improvement Yue et al. (2025). Although increasing the sampling number $k$ per prompt may seem like a straightforward solution, it typically yields only marginal improvements while significantly reducing sample efficiency Lin et al. (2025). This results in GRPO exhibiting limited exploration and low sample efficiency, and indirectly leads to training instability in the later stages due to the disappearance of reward signals Yu et al. (2025); Li et al. (2025).
To address the aforementioned issues, we propose EFRame: an Exploration-Filtering-Replay RL Framework. For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout phase to enable targeted exploration. During this process, the model generates a small number of high-quality positive samples along with a large number of low-quality negative samples. To reduce training variance and improve efficiency, we apply an online filtering mechanism to discard low-quality samples. Meanwhile, high-advantage samples are stored in a replay buffer and reused during training to amplify their impact and enhance model stability.
Previous studies have shown that both positive and negative samples contribute to RLVR training Zhu et al. (2025). With the support of the EFRame, we categorize training samples to better understand their roles in learning:
  • Regular samples refer to both correct and incorrect responses from prompts of moderate difficulty. They form the foundation of the reward signal, helping to improve model performance while ensuring training stability.
  • High-quality samples consist of correct responses obtained through exploration on difficult prompts. They are crucial for activating deeper model knowledge and enhancing the performance ceiling, but their effective use requires experience replay to avoid instability caused by their low sampling probability.
  • Low-quality samples refer to the numerous incorrect responses generated from challenging prompts. These samples offer minimal learning value and may even harm the training process by accelerating convergence to suboptimal solutions.
EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning. By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models. Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm. Moreover, by introducing a categorization of training samples based on their difficulty and informativeness, EFRame offers a new analysis on the roles of different sample within the RLVR framework. Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
2 Related Work
The early application of RL in LLMs and MLLMs OpenAI (2023); Team et al. (2024); Zhu et al. (2023); Wei et al. (2023); Liu et al. (2023) was primarily reflected in RLHF, where algorithms such as PPO were employed to align language model outputs with human preferences and Direct Preference Optimization (DPO) Rafailov et al. (2024) bypass the need for reward modeling and instead optimize the model policy directly from preference data. Recently, RL has advanced toward integrating rule-based reward mechanisms with sophisticated optimization techniques, as exemplified by models such as DeepSeek-R1 Guo et al. (2025) and Kimi-1.5 Team et al. (2025). As a lightweight variant of PPO, GRPO Shao et al. (2024) significantly improves training efficiency, but at the cost of increased instability. A series of subsequent studies have introduced optimizations from GRPO. These include gradient loss refinements, as demonstrated in DAPO Yu et al. (2025) and Dr.GRPO Liu et al. (2025); online filtering methods like Meng et al. (2025); advantage enhancement techniques, exemplified by VL-Rethinker Wang et al. (2025a) and Luffy Yan et al. (2025); clip acceleration methods such as CPPO Lin et al. (2025); curriculum learning strategies explored in Chen et al. (2025b; c); and unsupervised RL such as Wei et al. (2025a); Zhao et al. (2025); Agarwal et al. (2025). While existing methods typically focus on individual components, we systematically review the fundamental limitations of GRPO and propose a comprehensive framework with an integrated solution that addresses three critical aspects: exploration capacity, training stability, and sample efficiency. This helps create a more robust and effective GRPO-based training paradigm.
3 Overview of GRPO
3.1 Gradient Loss
GRPO has been employed to enhance the reasoning capabilities of large language models, particularly in mathematical problem solving. Compared to PPO, GRPO improves computational efficiency by eliminating the need for a separate critic model. Instead, it estimates the advantage directly using group-normalized rewards. Given a question $q$ and a corresponding verifiable answer $a$, GRPO generates a set of responses $O=\{o_{i}\}_{i=1}^{G}$ using the policy model $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{(q,a)\sim\mathcal{D},\ O=\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}(\cdot|q)}}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t},\ \operatorname{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]-\beta D_{KL}[\pi_{\theta}||\pi_{ref}]\right\},\end{split}$ (1)
where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i}<t)}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i}<t)}$ and $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}$.
3.2 Limitations of GRPO
Although GRPO has demonstrated strong performance in various settings, it encounters several practical challenges when applied to complex reasoning tasks. In the following discussion, we highlight three main issues: limited exploration capacity, training collapse, and low sample efficiency. These limitations reveal the need for more robust and adaptive RL strategies.
Limited Exploration Capacity. GRPO lacks effective mechanisms to promote exploration, leading to persistently low entropy levels in the early stages of training and an increased risk of convergence to local optima. Moreover, it does not differentiate between prompts of varying difficulty, which limits its ability to stimulate deeper reasoning capabilities. As a consequence, only superficial knowledge is activated, while deeper knowledge remains underutilized.
((a))
((b))
((c))
Figure 1: Increasing the sampling number to $G=20$ provides a performance improvement over $G=5$, but it incurs a fourfold increase in per-step runtime, leading to very low sample efficiency in practice.
Low Sample Efficiency. GRPO computes sample advantage based on relative advantage. When the number of rollouts is small, it suffers from a severe issue of advantage vanishing, leading to a lack of effective training signals. However, as the number of samples increases, the relative advantage becomes diluted, making it harder for high-quality samples to exert sufficient influence. This results in low sample efficiency: although the computational overhead grows nearly linearly with the number of samples, the corresponding performance improvement is significantly diminished, as shown in Fig. 1.
Instability. During the later stages of training, the GRPO algorithm may exhibit signs of instability or collapse, characterized by a sharp decline in reward and sudden spikes in both gradient and policy entropy, as illustrated in Fig. 2. In extreme cases, this can escalate into gradient explosion. Such instability presents a serious challenge to reliable long-horizon optimization, particularly in complex reasoning tasks where consistent gradient signals are critical for convergence. One contributing factor is the increased variance caused by the disappearance of relative advantages in the later stages of training. Supplementing with additional samples with non-zero advantages can help alleviate this issue.
((a))
((b))
((c))
Figure 2: When training Qwen2.5-VL-7B-Instruct on the Geo3k dataset using the GRPO algorithm, a collapse occurred.
4 Methodology
To address the issues in the previous section, we propose the Exploration-Filtering-Replay (EFR) framework, which consists of three main components: additional rollout, online filtering and replay buffer, as illustrated in Figure 3.
Figure 3: The overall workflow of EFR builds upon the vanilla GRPO framework by introducing three key components: additional rollout, online filtering, and replay buffer.
4.1 Regular Rollout
For each batch of prompts, we begin with a regular-rollout phase, following the original GRPO training process. This initial rollout serves two key purposes:
  • Obtain reward signals from regular (i.e., solvable) problems.
  • Identify more challenging prompts that fail to yield positive samples, which will be handled in subsequent stages.
Valuable samples are temporarily stored in the buffer for gradient updates.
4.2 Additional Rollout
GRPO computes loss over $G_{1}$ sampled outputs, making $G_{1}$ pivotal. Smaller $G_{1}$ limits learning capacity; while larger $G_{1}$ adds linear cost with limited benefit. For most questions, at least one positive sample can typically be found within a small number of rollouts. However, for more challenging problems, a larger number of samples and a higher temperature are often necessary to promote exploration and obtain meaningful reward signals.
To address this, we introduce an additional rollout mechanism: we perform an extra sampling round (e.g. $G_{2}$=20, temperature = 1.2) for challanging prompts.
Definition 4.1.
In GRPO-based algorithms with binary rewards (Premise, omitted in the following), difficult prompts are defined as those for which all sampled responses within the initial $G_{1}$ rollouts ($G_{1}$ is a smaller integer, such as 5) fail to obtain a positive reward, where $\bar{\mathcal{D}}\subseteq\mathcal{D}$ and:
$\bar{\mathcal{D}}=\left\{(q,a)~{}:~{}|\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}~{}:~{}\text{is\_equivalent}(a,o_{i})\right\}_{i=1}^{G_{1}}|=0\right\},$
The additional rollout is designed to activate deeper reasoning capabilities and increase the frequency of emergent “aha moments.”
4.3 Online filter
While additional rollout enhances the model’s ability to explore complex prompts, it also introduces computational overhead and potential training noise. In addition to filtering overlong samples, we design two additional filtering strategies to further enhance the stability and efficiency of the training process.
Regular Filter: In the first-round rollout, when the prompt is of regular difficulty—we adopt a conservative filtering strategy that retains all regular samples $O_{R}$ to balance efficiency and performance.
Definition 4.2.
Regular samples $O_{R}$ refer to any samples with a non-zero advantage on the prompt set $\mathcal{D}$:
$O_{R}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}}(\cdot|q):(q,a)\sim\mathcal{D},\ \hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{1}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{1}})}\neq 0\right\};$
At this stage, due to the moderate difficulty of prompts, all regular samples—both positive and negative—provide informative learning signals that contribute meaningfully to training.
Exploration Filter: When the prompt becomes challenging, we encourage the model to explore extensively in order to discover correct answers. Excluding samples with zero advantage, the remains can be categorized into the following two types.
Definition 4.3.
High-quality samples $O_{H}$ are positive samples for difficult prompts $\bar{\mathcal{D}}$:
$O_{H}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}>0\right\};$
low-quality samples are negative samples for difficult prompts $\bar{\mathcal{D}}$:
$O_{L}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}<0\right\}.$
Theorem 4.4.
For $o_{h}\in O_{H}$ and $o_{l}\in O_{L}$, $\sum\hat{A}_{h,0}+\sum\hat{A}_{l,0}=0$.
According to Definition 4.3 and Theorem 4.4, low-quality samples are significantly more numerous than high-quality ones, whereas the advantage values of high-quality samples are substantially higher. In such an exploration-driven setting, penalizing incorrect trajectories is not only inefficient but may also introduce noisy and unstable gradients. If gradients are computed from all samples, most of which are low quality, the informative signal from high-quality samples may be drowned out by chaotic updates from low-quality ones. To address this issue, we apply a more aggressive filtering strategy during the additional rollout—discarding all negative samples and retaining only positive samples for training. In practice, since this exploration phase targets only challenging prompts, the vast majority of generated samples (over 95%) are filtered out under this strategy.
To summarize, the policy gradient under our framework can be expressed as:
[…]
$\qquad\qquad\quad{}-\,\beta\,D_{\mathrm{KL}}\bigl{[}\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\bigr{]}\Bigr{)}\Biggr{]}.$
To prevent token-length bias, we compute the policy loss using the token-mean aggregation strategy.
4.4 Replay buffer
While GRPO is fundamentally an online RL algorithm, we find that experience replay with high-quality offline samples remains highly effective. In particular, trajectories discovered during the second-round rollout—often of higher quality due to deeper exploration and carry rich reward signals. By selectively storing high-quality samples in the replay buffer, we amplify their influence during training, particularly when their gradients are frequently clipped due to low sampling probabilities. Furthermore, by enabling the model to repeatedly leverage these rare yet informative samples, experience replay plays a pivotal role in enhancing stability of entropy.
Theorem 4.5 (Entropy Change under NPG Update).
If the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via natural policy gradient (Kakade, 2001) with step size $\eta$. Then the change in policy entropy under state between two steps approximately satisfies:
${H}(\pi_{\theta}^{k+1})-\displaystyle{H}(\pi_{\theta}^{k})\approx-\eta\cdot\mathbb{E}_{s\sim d_{\mu}^{k}}\text{Cov}_{a\sim\pi^{k}_{\theta}(\cdot\mid s)}\!\bigl{[}\log\!\pi^{k}_{\theta}(a\mid s),\,A^{\pi^{k}}(s,a)\bigr{]},$
This theorem was first introduced by Liu (2025), aand was organized and extended by cui2025entropymechanismreinforcementlearning. Proof can be seen in Liu (2025) and cui2025entropymechanismreinforcementlearning. ${H}$ indicates the policy entropy of policy model, and Cov denotes covariance, $\pi^{k}_{\theta}$ is the policy at step $k$, and $A^{\pi^{k}}(s,a)$ is the advantage function of action $a$ under state $s$. This result indicates that when high-advantage actions have low probability, they induce negative covariance, which increases entropy and reduce stability.
The high-quality samples identified through additional rollouts are typically responses to difficult prompts and thus are with relatively low probability. Using them only once yields limited and short-lived gradient effects, leading to an increase in model entropy. By incorporating these samples into the replay buffer, their probability of reoccurrence is effectively increased during training, enabling the model to transition from entropy increase to eventual convergence. In fact, the replay buffer here serves to reduce policy entropy; detailed discussions can be found in the following discussion section.
4.5 Discussion
The above subsections yield several practical insights regarding the interplay between Exploration-Filtering-Replay:
Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$. According Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing collapse of the entropy and maintaining effective exploration. At the same time, an excessive presence of low-quality samples can cause a nosedive of entropy, which is difficult to control. To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
Replay Buffer: By storing high-quality positive samples within the replay buffer, the model can repeatedly utilize their gradient information for subsequent updates. As the model has already been updated on these (problem, answer) pairs $(x,y)$, the recalculated $\log\pi_{\theta}(y\mid x)$ becomes larger upon rehearing these pairs. According to the theorem, this shift leads to a reduction in policy entropy, preventing excessive growth in entropy.
Applying any techniques in isolation may result in unhealthy dynamics of policy entropy. Subsequent ablation studies further corroborate these findings.
5 Experiments
To validate the effectiveness of our methods, we demonstrate our experiment setup and the results in the following sections.
Table 1: Comparative results across multiple multimodal reasoning benchmarks indicate that our method achieves excellent performance.
Models Size MathVision MathVerse MathVista We-Math Average
Close-Source Models
GPT-4o Hurst et al. (2024) - 33.95 48.83 59.50 65.00 51.82
Claude-3.5-Sonnet Anthropic (2024) - 46.48 57.64 68.20 73.05 61.34
Open-Source Multi-Modal Large Reasoning Models
QvQ-Preview Bai et al. (2025) 72B 35.56 52.81 69.20 65.29 55.72
Skywork R1V Peng et al. (2025) 38B 39.31 40.91 67.50 60.06 51.95
Open-Source 7B Multi-Modal Models
MM-Eureka Meng et al. (2025) 7B 28.36 49.52 70.40 66.03 53.58
OpenVLThinker Deng et al. (2025) 7B 25.30 42.79 64.10 63.91 49.03
ThinkLite-VL Wang et al. (2025b) 7B 27.71 47.41 69.00 67.99 53.03
VLAA-Thinker Chen et al. (2025a) 7B 26.61 47.49 66.20 66.49 51.70
Qwen2.5-VL-7B-Instruct with RL
Qwen2.5-VL Bai et al. (2025) 7B 24.87 43.83 66.30 62.87 49.47
Qwen2.5-VL+GRPO 7B 29.11 47.51 72.60 67.53 54.19
Qwen2.5-VL+DAPO 7B 27.92 48.48 72.30 69.08 54.45
Ours 7B 28.82 (+3.95) 48.81 (+4.98) 73.40 (+7.10) 69.48 (+6.61) 55.13 (+5.66)
5.1 Experimental setup
Training Datasets. Our experiments are conducted on two datasets to evaluate the effectiveness of the algorithm in both in-domain and out-of-domain settings: (1) Geometry3K focused on geometric problems Lu et al. (2021a), (2) a 6k out-of-domain dataset randomly sampled from the dataset proposed in  Wei et al. (2025b; a), which randomly sampled from established open-source resources, including Geometry3K Lu et al. (2021a), GeoQA Chen et al. (2021), GeoQA-Plus Cao & Xiao (2022), Geos Seo et al. (2015), AI2D Kembhavi et al. (2016), TQA Kim et al. (2018), FigureQA Kahou et al. (2017), TabMWP Lu et al. (2022b), ChartQA Masry et al. (2022), IconQA Lu et al. (2021b), Clevr-Math Lindström & Abraham (2022), M3CoT Chen et al. (2024), and ScienceQA Lu et al. (2022a).
Evaluation Benchmarks. Our experiments consist of two parts. First, we evaluate the overall performance of the algorithms on four open multimodal math benchmarks: MathVision Wang et al. (2024a), MathVista Lu et al. (2023), MathVerse Zhang et al. (2024), and We-Math Qiao et al. (2024). These benchmarks cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty. Second, we conduct a detailed analysis for algorithms on the Geometry3K dataset to examine the algorithmic behaviors.
Baselines. We compare our models with a range of advanced multimodal large language models. These include GPT-4o Hurst et al. (2024) and Claude-3.5-Sonnet Anthropic (2024), which are state-of-the-art closed-source MLLMs. QvQ-Preview-72B[2] and Skywork R1V-32B  Peng et al. (2025) are multimodal large reasoning models specifically designed for R1-like chain-of-thought capabilities. At the 7B scale, we compare with MM-Eureka-7B Meng et al. (2025), OpenVLThinker-7B Deng et al. (2025), ThinkLite-VL-7B Wang et al. (2025b), and VLAA-Thinker-Qwen2.5VL-7B Chen et al. (2025a). All of these 7B models are based on Qwen2.5-VL and are enhanced for mathematical reasoning via fine-tuning, reinforcement learning, or iterative self-improvement techniques. In addition, we compare the performance of GRPO and DAPO on the Qwen2.5-VL-7B-Instruct model to assess the impact of different strategies under consistent model and training settings.
Implementation details. Our work is based on EasyR1 and VeRL framework Yaowei et al. (2025); Sheng et al. (2025). We initialize all policy models with Qwen2.5-VL-7B-Instruct and conduct experiments on 8 A800 GPUs (40GB). Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout $G_{1}$=5, temperature 1.0, and learning rate $10^{-6}$. GRPO and our method adopt $\epsilon_{low}=\epsilon_{high}=0.2$ and a binary reward to avoid relative advantage bias (see detail in Appendix A), while DAPO includes $\epsilon_{high}=0.3$ and an additional length-aware linear penalty (up to 0.5) for responses exceeding 1536 tokens. For our method, we perform additional rollouts with $G_{2}=20$, temperature = 1.2; responses with absolute advantage larger than 1 are stored in a replay buffer of size 5120, with experience replay triggered every 5 steps.
Table 2: The performance comparison on the Geometry3K dataset shows that our method achieves a significant improvement over existing approaches.
Models Qwen2.5-VL-Instruct Ours GRPO DAPO
Accuracy 38.44 55.24(+16.80) 50.75 (+12.31) 49.09 (+10.65)
Time/step(s) - 442($\times$1.35) 327($\times$1.00) 182($\times$0.57)
5.2 Main results
((a))
((b))
((c))
Figure 4: On the Geometry3K dataset, our method not only achieves excellent performance but also demonstrates superior exploration capability and training stability.
Performance. In our experiments, we employ the Exploration-Filtering-Replay framework, yielding notable improvements on both out-of-domain and in-domain datasets. As shown in Table 1, the proposed method delivers consistently strong results across four multimodal reasoning benchmarks, with an average gain of +5.66 over the base model (Qwen2.5-VL-7B-Instruct), outperforming GRPO by 19.9%. On the in-domain Geometry3K dataset, the improvement is even more pronounced: Table 2 reports a +16.80 gain, representing a 36.5% relative advantage over GRPO.
Exploration. As shown in Fig. 4, our framework maintains the policy entropy around 1, striking a balance between sufficient exploration and exploitation. In the later training stages, when the vanilla GRPO tends to suffer from vanishing advantages, our method further increases entropy to encourage continued exploration of informative reward signals.
Stability. Our framework leverages additional rollouts to enrich reward signals, while an online filtering mechanism mitigates training variance by excluding low-quality samples. Extensive experiments reveal that the proposed method circumvents collapse issues—occasionally encountered in GRPO and DAPO. Furthermore, the steadily increasing response length during training reflects the model’s progressively improved reasoning ability.
6 Ablation Studies
To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and replay buffer—are all essential for achieving the observed improvements in stability and overall performance.
6.1 Additional Rollout
Disabling the additional rollout module exposes the model to a fundamental trade-off inherent in GRPO-based training. When the number of samples $G_{1}$ is small, the reward signal is constrained by the base model’s pass@$G_{1}$, limiting the model’s ability to solve challenging problems and suppressing the emergence of “aha moments.”
Claim 1.
High-quality samples have the potential to activate deeper reasoning abilities within the model.
As shown in Fig. 5, the model’s reward on the training set plateaus at a relatively low level, restricting further gains in test accuracy. In the early stages of training, the entropy remains low, indicating limited exploration. However, once the model hits this performance bottleneck, entropy begins to rise—yet this delayed exploration yields minimal returns, suggesting that the model struggles to escape local optima. The additional rollout module is specifically designed to break this situation, enabling targeted deep exploration.
((a))
((b))
((c))
Figure 5: The ablation study conducted on the Geometry3K dataset not only validates the effectiveness of our proposed method but also reveals the distinct roles played by samples of varying quality during the training process.
6.2 Online Filtering
Removing the online filtering module results in significant instability and inefficiency during training. Without filtering, the model treats all samples—regardless of their advantage—as equally important, allowing zero-advantage samples to increase gradient variance and low-quality negative trajectories to distorb the learning signal. This problem is further exacerbated during the additional rollout phase, where abundant incorrect samples significantly amplifies gradient variance, making it difficult for the model to retain meaningful patterns from rare high-quality responses. As a result, training becomes slower, more chaotic, and more prone to divergence.
Claim 2.
Low-quality samples from challenging prompts often mislead optimization, causing the model to converge prematurely to suboptimal solutions and impeding further learning.
As shown in Fig. 5, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set, but yields little improvement on the test set. This indicates that most low-quality samples lead the model to converge more quickly to a local optimum, offering little benefit to its generalization ability. Moreover, the forward and backward computation of these negative samples incurs substantial overhead, reducing the sample efficiency of the overall training process. Online filtering plays a crucial role in maintaining gradient clarity and training stability by selectively preserving informative samples and discarding neutral or low-impact ones.
6.3 Replay Buffer
When the replay buffer is disabled, we observe a drastic drop in performance, even falling below that of the vanilla GRPO baseline, as shown in Fig. 5. This counterintuitive result reveals that the abundant high-quality samples generated during the additional rollout phase can actually become detrimental without a proper replay mechanism.
Claim 3.
While high-quality samples may induce high-entropy exploration, incorporating experience replay enables the model to retain and consolidate useful gradients, thus restoring training stability.
As high-quality samples typically have low generation probabilities, rewarding them shifts the model’s response distribution toward low-probability but correct outputs, disrupting the convergence trajectory and consequently increasing entropy. Due to their high advantage values and low sampling probabilities, high-quality samples are more likely to have their gradients clipped during training. In the absence of a replay buffer, such valuable samples contribute only short-lived and unstable gradients that are quickly overwritten. Instead of reinforcing beneficial behavior, this transient gradient signal increases policy uncertainty, pushing the model into a state of chaotic, high-entropy exploration, ultimately harming both learning stability and performance. To address this issue, we leverage experience replay to ensure that high-quality samples continuously provide effective gradients, enabling the model to converge along the new trajectory.
7 Conclusion and Discussion
In this work, we investigate the impact of sample quality on RL for large language models, with a particular focus on addressing challenging prompts. Our findings show that low-quality incorrect samples—those with low absolute advantage—tend to trap the model in local optima and offer little or even negative contribution to training. In contrast, high-quality samples with large advantages play a critical role in guiding the model through a meaningful transition from entropy-driven exploration to entropy-reducing convergence. This process is essential for eliciting deeper reasoning capabilities within the model. To this end, we propose an Exploration-Filtering-Replay framework that emphasizes the discovery, selection, and sustained utilization of high-quality samples. Our method introduces additional rollouts to enhance reward density, employs online filtering to suppress noisy samples, and integrates experience replay to amplify the impact of high-quality samples across training iterations. This strategy not only helps transform isolated “aha moments” into a sustained “aha state,” but also shifts reinforcement learning from a “reinforcement lottery” toward a more structured and reliable process of reinforcement learning.
Appendix
Figure 6: During training with GRPO, when the primary reward signal (accuracy) vanishes, other auxiliary rewards become dominant, which introduces bias.
Appendix A Relative Advantage Bias of GRPO
In GRPO, computing relative advantages within a group can introduce bias, especially when the reward becomes multi-dimensional. For example, as shown in Fig. 6, if the reward comprises both accuracy and auxiliary signals such as length, format, or entropy, a significant issue arises when the accuracy signal vanishes within a group—i.e., when all responses are either correct or incorrect. In such cases, auxiliary rewards (e.g., length) dominate the advantage computation. This may result in incorrect responses being rewarded simply for being shorter, or correct responses being penalized for being longer. Such misleading gradients introduce irreversible noise into training and lead to reward hacking. Therefore, in the training involved in this paper, we use binary rewards whenever possible.
Appendix B Key Hyperparameters
The quantity of high-quality samples varies significantly with dataset difficulty, which makes our algorithm sensitive to three key hyperparameters: clip higher, replay buffer size, and replay frequency. This is because the gradients from high-quality samples must be just perfect—between underfitting, proper fitting, and overfitting—to be effective. Notably, on relatively simple datasets such as Math, where high-quality samples are inherently scarce, the improvements of our framework over the original GRPO are less pronounced.

### v2

@@
-EFRame: Deeper Reasoning via Exploration-Filtering-Replay Reinforcement Learning Framework
+EFRame: Deeper Reasoning via Exploration-Filter-Replay Reinforcement Learning Framework
 Abstract
@@
 Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks.
-To address these limitations, we introduce EFRame, an Exploration-Filtering-Replay framework that systematically augments GRPO along three critical dimensions.
+To address these limitations, we introduce EFRame, an Exploration-Filter-Replay framework that systematically augments GRPO along three critical dimensions.
 EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples.
@@
 More recently, the focus has shifted towards Reinforcement Learning with Verifiable Rewards (RLVR), which seeks to provide more consistent and scalable supervision by leveraging structured, verifiable reward signals Mroueh (2025).
-Within this paradigm, DeepSeek-R1 introduced Group Relative Policy Optimization (GRPO), a lightweight and efficient variant of PPO that compute relative rewards within rollout groups to eliminate the need for the critic model Shao et al.
+Within this paradigm, DeepSeek-R1 introduced Group Relative Policy Optimization (GRPO), a lightweight and efficient variant of PPO that computes relative rewards within rollout groups to eliminate the need for the critic model Shao et al.
 (2024); Liu et al.
@@
 (2025).
-This results in GRPO exhibiting limited exploration and low sample efficiency, and indirectly leads to training instability in the later stages due to the disappearance of reward signals Yu et al.
+This results in GRPO exhibiting limited exploration, and low sample efficiency, and indirectly leads to training instability in the later stages due to the disappearance of reward signals Yu et al.
 (2025); Li et al.
 (2025).
-To address the aforementioned issues, we propose EFRame: an Exploration-Filtering-Replay RL Framework.
+To address the aforementioned issues, we propose EFRame: an Exploration-Filter-Replay RL Framework.
 For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout phase to enable targeted exploration.
@@
 (2025).
-With the support of the EFRame, we categorize training samples to better understand their roles in learning:
+With the support of the EFRame, we categorize training samples to understand their roles in learning better:
   • Regular samples refer to both correct and incorrect responses from prompts of moderate difficulty.
 They form the foundation of the reward signal, helping to improve model performance while ensuring training stability.
-  • High-quality samples consist of correct responses obtained through exploration on difficult prompts.
+  • High-quality samples consist of correct responses obtained through the exploration of difficult prompts.
 They are crucial for activating deeper model knowledge and enhancing the performance ceiling, but their effective use requires experience replay to avoid instability caused by their low sampling probability.
@@
 Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm.
-Moreover, by introducing a categorization of training samples based on their difficulty and informativeness, EFRame offers a new analysis on the roles of different sample within the RLVR framework.
+Moreover, by introducing a categorization of training samples based on their difficulty and informativeness, EFRame offers a new analysis of the roles of different samples within the RLVR framework.
 Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
@@
 (2025); curriculum learning strategies explored in Chen et al.
-(2025b; c); and unsupervised RL such as Wei et al.
+(2025b; c); and unsupervised RL such as Zhang et al.
+(2025); Wei et al.
 (2025a); Zhao et al.
@@
 3.1 Gradient Loss
-GRPO has been employed to enhance the reasoning capabilities of large language models, particularly in mathematical problem solving.
+GRPO has been employed to enhance the reasoning capabilities of large language models, particularly in mathematical problem-solving.
 Compared to PPO, GRPO improves computational efficiency by eliminating the need for a separate critic model.
@@
 2.
-In extreme cases, this can escalate into gradient explosion.
+In extreme cases, this can escalate into a gradient explosion.
 Such instability presents a serious challenge to reliable long-horizon optimization, particularly in complex reasoning tasks where consistent gradient signals are critical for convergence.
@@
 4 Methodology
-To address the issues in the previous section, we propose the Exploration-Filtering-Replay (EFR) framework, which consists of three main components: additional rollout, online filtering and replay buffer, as illustrated in Figure 3.
-Figure 3: The overall workflow of EFR builds upon the vanilla GRPO framework by introducing three key components: additional rollout, online filtering, and replay buffer.
+To address the issues in the previous section, we propose the EFRame, which consists of three main components: additional rollout, online filtering and replay buffer, as illustrated in Figure 3.
+Figure 3: The overall workflow of EFRame builds upon the vanilla GRPO framework by introducing three key components: additional rollout, online filtering, and replay buffer.
 4.1 Regular Rollout
-For each batch of prompts, we begin with a regular-rollout phase, following the original GRPO training process. This initial rollout serves two key purposes:
+For each batch of prompts, we begin with a regular rollout phase, following the original GRPO training process. This initial rollout serves two key purposes:
   • Obtain reward signals from regular (i.e., solvable) problems.
@@
 However, for more challenging problems, a larger number of samples and a higher temperature are often necessary to promote exploration and obtain meaningful reward signals.
-To address this, we introduce an additional rollout mechanism: we perform an extra sampling round (e.g. $G_{2}$=20, temperature = 1.2) for challanging prompts.
+To address this, we introduce an additional rollout mechanism: we perform an extra sampling round (e.g. $G_{2}$=20, temperature = 1.2) for challenging prompts.
 Definition 4.1.
@@
 While GRPO is fundamentally an online RL algorithm, we find that experience replay with high-quality offline samples remains highly effective.
-In particular, trajectories discovered during the second-round rollout—often of higher quality due to deeper exploration and carry rich reward signals.
+In particular, trajectories discovered during the second-round rollout—are often of higher quality due to deeper exploration and carry rich reward signals.
 By selectively storing high-quality samples in the replay buffer, we amplify their influence during training, particularly when their gradients are frequently clipped due to low sampling probabilities.
-Furthermore, by enabling the model to repeatedly leverage these rare yet informative samples, experience replay plays a pivotal role in enhancing stability of entropy.
+Furthermore, by enabling the model to repeatedly leverage these rare yet informative samples, experience replay plays a pivotal role in enhancing the stability of entropy.
 Theorem 4.5 (Entropy Change under NPG Update).
@@
 ${H}(\pi_{\theta}^{k+1})-\displaystyle{H}(\pi_{\theta}^{k})\approx-\eta\cdot\mathbb{E}_{s\sim d_{\mu}^{k}}\text{Cov}_{a\sim\pi^{k}_{\theta}(\cdot\mid s)}\!\bigl{[}\log\!\pi^{k}_{\theta}(a\mid s),\,A^{\pi^{k}}(s,a)\bigr{]},$
-This theorem was first introduced by Liu (2025), aand was organized and extended by cui2025entropymechanismreinforcementlearning.
-Proof can be seen in Liu (2025) and cui2025entropymechanismreinforcementlearning.
+This theorem was first introduced by Liu (2025), and was organized and extended by Cui et al.
+(2025).
+Proof can be seen in Liu (2025) and Cui et al.
+(2025).
 ${H}$ indicates the policy entropy of policy model, and Cov denotes covariance, $\pi^{k}_{\theta}$ is the policy at step $k$, and $A^{\pi^{k}}(s,a)$ is the advantage function of action $a$ under state $s$.
-This result indicates that when high-advantage actions have low probability, they induce negative covariance, which increases entropy and reduce stability.
+This result indicates that when high-advantage actions have low probability, they induce negative covariance, which increases entropy and reduces stability.
 The high-quality samples identified through additional rollouts are typically responses to difficult prompts and thus are with relatively low probability.
@@
 4.5 Discussion
-The above subsections yield several practical insights regarding the interplay between Exploration-Filtering-Replay:
-Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$.
-According Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing collapse of the entropy and maintaining effective exploration.
+The above subsections yield several practical insights regarding the interplay between Exploration-Filter-Replay:
+Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high-temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$.
+According to Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing the collapse of the entropy and maintaining effective exploration.
 At the same time, an excessive presence of low-quality samples can cause a nosedive of entropy, which is difficult to control.
@@
 These benchmarks cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty.
-Second, we conduct a detailed analysis for algorithms on the Geometry3K dataset to examine the algorithmic behaviors.
+Second, we conduct a detailed analysis of algorithms on the Geometry3K dataset to examine the algorithmic behaviors.
 Baselines.
@@
 Performance.
-In our experiments, we employ the Exploration-Filtering-Replay framework, yielding notable improvements on both out-of-domain and in-domain datasets.
+In our experiments, we employ the Exploration-Filter-Replay framework, yielding notable improvements on both out-of-domain and in-domain datasets.
 As shown in Table 1, the proposed method delivers consistently strong results across four multimodal reasoning benchmarks, with an average gain of +5.66 over the base model (Qwen2.5-VL-7B-Instruct), outperforming GRPO by 19.9%.
@@
 Removing the online filtering module results in significant instability and inefficiency during training.
-Without filtering, the model treats all samples—regardless of their advantage—as equally important, allowing zero-advantage samples to increase gradient variance and low-quality negative trajectories to distorb the learning signal.
-This problem is further exacerbated during the additional rollout phase, where abundant incorrect samples significantly amplifies gradient variance, making it difficult for the model to retain meaningful patterns from rare high-quality responses.
+Without filtering, the model treats all samples—regardless of their advantage—as equally important, allowing zero-advantage samples to increase gradient variance and low-quality negative trajectories to distort the learning signal.
+This problem is further exacerbated during the additional rollout phase, where abundant incorrect samples significantly amplify gradient variance, making it difficult for the model to retain meaningful patterns from rare high-quality responses.
 As a result, training becomes slower, more chaotic, and more prone to divergence.
@@
 As shown in Fig.
-5, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set, but yields little improvement on the test set.
+5, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set but yields little improvement on the test set.
 This indicates that most low-quality samples lead the model to converge more quickly to a local optimum, offering little benefit to its generalization ability.
@@
 While high-quality samples may induce high-entropy exploration, incorporating experience replay enables the model to retain and consolidate useful gradients, thus restoring training stability.
-As high-quality samples typically have low generation probabilities, rewarding them shifts the model’s response distribution toward low-probability but correct outputs, disrupting the convergence trajectory and consequently increasing entropy.
+As high-quality samples typically have low generation probabilities, rewarding them shifts the model’s response distribution toward low probability but correct outputs, disrupting the convergence trajectory and consequently increasing entropy.
 Due to their high advantage values and low sampling probabilities, high-quality samples are more likely to have their gradients clipped during training.
@@
 This process is essential for eliciting deeper reasoning capabilities within the model.
-To this end, we propose an Exploration-Filtering-Replay framework that emphasizes the discovery, selection, and sustained utilization of high-quality samples.
+To this end, we propose an Exploration-Filter-Replay framework that emphasizes the discovery, selection, and sustained utilization of high-quality samples.
 Our method introduces additional rollouts to enhance reward density, employs online filtering to suppress noisy samples, and integrates experience replay to amplify the impact of high-quality samples across training iterations.
@@
 Appendix
-Figure 6: During training with GRPO, when the primary reward signal (accuracy) vanishes, other auxiliary rewards become dominant, which introduces bias.
 Appendix A Relative Advantage Bias of GRPO
@@
 For example, as shown in Fig.
-6, if the reward comprises both accuracy and auxiliary signals such as length, format, or entropy, a significant issue arises when the accuracy signal vanishes within a group—i.e., when all responses are either correct or incorrect.
+6, if the reward comprises both accuracy and auxiliary signals such as length, format, or entropy, a significant issue arises when the accuracy signal vanishes within a group—i.e.
+when all responses are either correct or incorrect.
 In such cases, auxiliary rewards (e.g., length) dominate the advantage computation.
@@
 Therefore, in the training involved in this paper, we use binary rewards whenever possible.
+Figure 6: During training with GRPO, when the primary reward signal (accuracy) vanishes, other auxiliary rewards become dominant, which introduces bias.
 Appendix B Key Hyperparameters

### v3

@@
 Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO.
-Furthermore, EFRame enables a more fine-grained categorization of training samples, allowing for a deeper analysis of how different types of samples contribute to the learning process in RL.
+Furthermore, EFRame not only enables fine-grained categorization of training samples for deeper insight into their contributions, but also introduces an efficient and precise mechanism for entropy control, which is critical for balancing exploration and convergence in RL training.
 Our code is available at https://github.com/597358816/EFRame.
@@
 Meanwhile, high-advantage samples are stored in a replay buffer and reused during training to amplify their impact and enhance model stability.
-Previous studies have shown that both positive and negative samples contribute to RLVR training Zhu et al.
-(2025).
-With the support of the EFRame, we categorize training samples to understand their roles in learning better:
-  • Regular samples refer to both correct and incorrect responses from prompts of moderate difficulty.
-They form the foundation of the reward signal, helping to improve model performance while ensuring training stability.
-  • High-quality samples consist of correct responses obtained through the exploration of difficult prompts.
-They are crucial for activating deeper model knowledge and enhancing the performance ceiling, but their effective use requires experience replay to avoid instability caused by their low sampling probability.
-  • Low-quality samples refer to the numerous incorrect responses generated from challenging prompts.
-These samples offer minimal learning value and may even harm the training process by accelerating convergence to suboptimal solutions.
+The contributions of EFRame are threefold:
+  1. 1.
+    Enhanced High-Difficulty Reasoning: EFRame significantly improves the performance of GRPO on challenging reasoning tasks.
+By integrating exploration, filtering, and replay into a unified framework, it enables more complete and deeper reasoning processes in large language models.
+  2. 2.
+    Sample-Type-Aware Training: EFRame introduces a principled categorization of training samples to better understand their distinct roles in learning.
+It demonstrates that high-advantage, low-probability positive samples are crucial for triggering exploration and unlocking deeper knowledge, whereas low-advantage negative samples tend to drive premature convergence to suboptimal solutions and should be suppressed.
+  3. 3.
+    Precise and Efficient Entropy Control: EFRame offers a practical and effective approach to entropy control by jointly tuning the additional rollout temperature and the number of replayed samples.
+This mechanism provides more flexible and stable entropy regulation than traditional reward-based entropy regularization, supporting both efficient exploration and stable convergence.
 EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning.
@@
 Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm.
-Moreover, by introducing a categorization of training samples based on their difficulty and informativeness, EFRame offers a new analysis of the roles of different samples within the RLVR framework.
 Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
@@
 In addition to filtering overlong samples, we design two additional filtering strategies to further enhance the stability and efficiency of the training process.
-Regular Filter: In the first-round rollout, when the prompt is of regular difficulty—we adopt a conservative filtering strategy that retains all regular samples $O_{R}$ to balance efficiency and performance.
+Regular Filter: In the first-round rollout, when the prompt is of regular difficulty, we adopt a conservative filtering strategy that retains all regular samples $O_{R}$ to balance efficiency and performance.
 Definition 4.2.
@@
 If the actor policy $\pi_{\theta}$ is a tabular softmax policy updated via natural policy gradient (Kakade, 2001) with step size $\eta$.
-Then the change in policy entropy under state between two steps approximately satisfies:
+Then the change in policy entropy between two steps approximately satisfies:
 ${H}(\pi_{\theta}^{k+1})-\displaystyle{H}(\pi_{\theta}^{k})\approx-\eta\cdot\mathbb{E}_{s\sim d_{\mu}^{k}}\text{Cov}_{a\sim\pi^{k}_{\theta}(\cdot\mid s)}\!\bigl{[}\log\!\pi^{k}_{\theta}(a\mid s),\,A^{\pi^{k}}(s,a)\bigr{]},$
@@
 According to Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing the collapse of the entropy and maintaining effective exploration.
-At the same time, an excessive presence of low-quality samples can cause a nosedive of entropy, which is difficult to control.
+At the same time, an excessive presence of low-quality samples can cause a nosedive in entropy, which is difficult to control.
 To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
@@
 According to the theorem, this shift leads to a reduction in policy entropy, preventing excessive growth in entropy.
-Applying any techniques in isolation may result in unhealthy dynamics of policy entropy. Subsequent ablation studies further corroborate these findings.
 5 Experiments
@@
 Implementation details.
-Our work is based on EasyR1 and VeRL framework Yaowei et al.
+Our work is based on EasyR1 and VeRL frameworks Yaowei et al.
 (2025); Sheng et al.
@@
 Furthermore, the steadily increasing response length during training reflects the model’s progressively improved reasoning ability.
+5.3 Entropy Control
+To achieve a well-balanced exploration process, we incorporate entropy control into EFRame by adjusting two key hyperparameters: the additional rollout temperature $t_{a}$ and the replay buffer size $R_{s}$.
+Specifically, increasing the temperature in the additional rollout phase encourages broader exploration, enabling the model to discover low-probability but potentially high-reward trajectories.
+At the same time, we increase the $R_{s}$ to promote convergence and prevent excessive entropy growth.
+Together, these mechanisms allow for a controlled yet sufficiently rich exploration process—ensuring that the model explores meaningfully without destabilizing training due to overly random behavior.
+((a))
+((b))
+((c))
+((d))
+((e))
+((f))
+Accuracy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
+$R_{s}=5120$ 53.74 52.41 55.24
+$R_{s}=2560$ 52.75 54.41 52.75
+$R_{s}=0$ 52.41 51.08 50.25
+Entropy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
+$R_{s}=5120$ 0.089 0.305 0.726
+$R_{s}=2560$ 0.091 0.312 2.366
+$R_{s}=0$ 0.092 2.437 4.288
+Figure 5: We report the peak accuracy within the first 200 training steps and the average entropy over the first 150 steps.
+Experiments on the GEO3K dataset using Qwen2.5-VL-7B-Instruct under different settings of additional rollout temperature $t_{a}$ and replay buffer size $R_{s}$ demonstrate that increasing $t_{a}$ enhances exploration, reflected in higher entropy, while enlarging $R_{s}$ promotes convergence, leading to lower entropy.
+Optimal performance is achieved when exploration and convergence are well-balanced.
+We observe in Fig.
+5 that increasing the additional rollout temperature encourages broader exploration, leading to a rise in policy entropy.
+In contrast, enlarging the replay buffer size, which increases the number of high-quality samples reused during training, promotes convergence and results in decreased entropy.
+These two hyperparameters exhibit opposing effects on entropy dynamics, providing a practical mechanism for entropy control within EFRame.
+We find that model performance peaks when the entropy stabilizes around a moderate level.
+Fig.
+5 reports the accuracy under different $t_{a}$ and $R_{s}$, highlighting the impact of entropy control of EFRame on learning efficacy.
+Our experiments demonstrate that entropy control can be achieved with high precision by combining high-temperature positive sampling and experience replay, which are more effective than using entropy-based reward functions.
 6 Ablation Studies
-To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and replay buffer—are all essential for achieving the observed improvements in stability and overall performance.
+To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and replay buffer—are essential for achieving the observed improvements in stability and overall performance.
 6.1 Additional Rollout
@@
 As shown in Fig.
-5, the model’s reward on the training set plateaus at a relatively low level, restricting further gains in test accuracy.
+6, the model’s reward on the training set plateaus at a relatively low level, restricting further gains in test accuracy.
 In the early stages of training, the entropy remains low, indicating limited exploration.
@@
 As shown in Fig.
-5, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set but yields little improvement on the test set.
+6, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set but yields little improvement on the test set.
 This indicates that most low-quality samples lead the model to converge more quickly to a local optimum, offering little benefit to its generalization ability.
@@
 When the replay buffer is disabled, we observe a drastic drop in performance, even falling below that of the vanilla GRPO baseline, as shown in Fig.
-5.
+6.
 This counterintuitive result reveals that the abundant high-quality samples generated during the additional rollout phase can actually become detrimental without a proper replay mechanism.
@@
 Our method introduces additional rollouts to enhance reward density, employs online filtering to suppress noisy samples, and integrates experience replay to amplify the impact of high-quality samples across training iterations.
-This strategy not only helps transform isolated “aha moments” into a sustained “aha state,” but also shifts reinforcement learning from a “reinforcement lottery” toward a more structured and reliable process of reinforcement learning.
+This strategy not only helps transform isolated “aha moments” into a sustained “aha state” but also shifts reinforcement learning from a “reinforcement lottery” toward a more structured and reliable process of reinforcement learning.
 Appendix
@@
 For example, as shown in Fig.
-6, if the reward comprises both accuracy and auxiliary signals such as length, format, or entropy, a significant issue arises when the accuracy signal vanishes within a group—i.e.
+7, if the reward comprises both accuracy and auxiliary signals such as length, format, or entropy, a significant issue arises when the accuracy signal vanishes within a group—i.e.
 when all responses are either correct or incorrect.
@@
 Figure 7: During training with GRPO, when the primary reward signal (accuracy) vanishes, other auxiliary rewards become dominant, which introduces bias.
-Appendix B Key Hyperparameters
-The quantity of high-quality samples varies significantly with dataset difficulty, which makes our algorithm sensitive to three key hyperparameters: clip higher, replay buffer size, and replay frequency.
-This is because the gradients from high-quality samples must be just perfect—between underfitting, proper fitting, and overfitting—to be effective.
-Notably, on relatively simple datasets such as Math, where high-quality samples are inherently scarce, the improvements of our framework over the original GRPO are less pronounced.

### v5

@@
-EFRame: Deeper Reasoning via Exploration-Filter-Replay Reinforcement Learning Framework
+inkscapearea=page
 Abstract
 Recent advances in reinforcement learning (RL) have significantly enhanced the reasoning capabilities of large language models (LLMs).
-Group Relative Policy Optimization (GRPO), an efficient variant of PPO that lowers RL’s computational cost, still faces limited exploration, low sample efficiency and instability, constraining its performance on complex reasoning tasks.
-To address these limitations, we introduce EFRame, an Exploration-Filter-Replay framework that systematically augments GRPO along three critical dimensions.
-EFRame performs additional rollouts to explore high-quality trajectories, applies online filtering to eliminate low-quality samples that introduce noise and variance, and leverages experience replay to repeatedly exploit rare but informative samples.
-EFRame establishes a complete and stable learning cycle, guiding the model through a structured transition from exploration to convergence.
-Our experiments across a variety of reasoning benchmarks demonstrate that EFRame not only improves the robustness and efficiency of training, but also enables access to deeper reasoning capabilities that remain unattainable under vanilla GRPO.
-Furthermore, EFRame not only enables fine-grained categorization of training samples for deeper insight into their contributions, but also introduces an efficient and precise mechanism for entropy control, which is critical for balancing exploration and convergence in RL training.
-Our code is available at https://github.com/597358816/EFRame.
+Group Relative Policy Optimization (GRPO), a lightweight variant of Proximal Policy Optimization (PPO), improves efficiency but suffers from limited exploration and training instability, limiting its effectiveness on complex reasoning tasks.
+To address these challenges, we introduce EFRame, an Exploration-Filter-Replay framework that augments GRPO across three dimensions: additional rollouts enable deeper and more targeted exploration, online filtering removes low-quality samples to stabilize gradients and accelerate training, and experience replay amplifies rare yet informative trajectories for stable convergence.
+This unified framework establishes a principled training cycle that balances exploration, efficiency, and stability.
+Experiments on diverse reasoning benchmarks demonstrate that EFRame achieves consistent gains, including a 37.9% relative improvement on Geometry3K over GRPO.
+EFRame further supports fine-grained sample categorization and precise entropy control, highlighting it as a robust solution for advancing deeper reasoning in LLMs.
 1 Introduction
@@
 GRPO has demonstrated remarkable improvements in training efficiency and has shown promising potential in enhancing the reasoning capabilities of LLMs.
-RLVR relies on extracting reward signals from the model’s own pass@$k$ outputs to improve its pass@1 performance.
-However, this design inherently couples final performance to the choice of $k$, as the reward signal vanishes when all sampled responses are either correct or incorrect, making it impossible for the model to extract useful knowledge from such data.
-This constrains RLVR’s learning capacity to the base model’s pass@$k$, limiting further exploration and improvement Yue et al.
+Despite its efficiency, GRPO suffers from limited exploration and training instability on complex reasoning tasks.
+We conducted extensive experiments and categorized the failed GRPO training cases into two types, as illustrated in Fig.
+1, which shows representative runs on Qwen2.5-VL-7B-Instruct with the Geometry3K dataset.
+In the first example (GRPO-1 in Fig.
+1), exploration was severely limited: the policy entropy continuously decreased from an initial value of 0.25 throughout training, making it difficult for the model to perform effective exploration and resulting in insufficient fine-tuning improvements.
+In the second example (GRPO-2 in Fig.
+1), the training suffered an irreversible collapse, marked by an arch-shaped fall of rewards and accuracy during mid-training, followed by a breakdown of model performance that could not be recovered.
+((a))
+((b))
+((c))
+Figure 1: Two prominent issues of GRPO: limited exploration and training instability, when training on Qwen2.5-VL-7B-Instruct with the Geometry3K dataset.
+GRPO derives reward signals from relative advantages computed within rollout groups.
+This design inherently ties the training process to the accuracy of groups Yue et al.
 (2025).
-Although increasing the sampling number $k$ per prompt may seem like a straightforward solution, it typically yields only marginal improvements while significantly reducing sample efficiency Lin et al.
-(2025).
-This results in GRPO exhibiting limited exploration, and low sample efficiency, and indirectly leads to training instability in the later stages due to the disappearance of reward signals Yu et al.
+When all sampled responses are either correct or incorrect, the relative advantage collapses to zero, yielding no effective learning signal.
+In long-horizon training, this advantage-vanishing effect amplifies gradient variance, which often manifests as gradient explosions and continuous reward degradation, eventually leading to irreversible training collapse Yu et al.
 (2025); Li et al.
 (2025).
+Collectively, these issues constrain GRPO’s ability to sustain exploration, preserve model accuracy, and maintain stability on challenging reasoning tasks.
 To address the aforementioned issues, we propose EFRame: an Exploration-Filter-Replay RL Framework.
-For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout phase to enable targeted exploration.
+For challenging prompts that cannot be effectively handled by regular rollouts, EFRame introduces an additional rollout to enable targeted exploration.
 During this process, the model generates a small number of high-quality positive samples along with a large number of low-quality negative samples.
 To reduce training variance and improve efficiency, we apply an online filtering mechanism to discard low-quality samples.
-Meanwhile, high-advantage samples are stored in a replay buffer and reused during training to amplify their impact and enhance model stability.
+Meanwhile, high-advantage samples are stored for experience replay to amplify their impact and enhance model stability.
 The contributions of EFRame are threefold:
@@
   3. 3.
-    Precise and Efficient Entropy Control: EFRame offers a practical and effective approach to entropy control by jointly tuning the additional rollout temperature and the number of replayed samples.
-This mechanism provides more flexible and stable entropy regulation than traditional reward-based entropy regularization, supporting both efficient exploration and stable convergence.
-EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in reinforcement learning.
+    Efficient Entropy Control: EFRame provides a practical and effective method for entropy control by jointly tuning the additional rollout temperature and the number of replayed samples.
+Compared to traditional reward-based entropy regularization, this mechanism enables more flexible and stable entropy regulation, thereby facilitating both efficient exploration and stable convergence.
+Overall, EFRame provides a systematic solution to the inherent trade-offs among stability, efficiency, and exploration capacity in RL.
+By effectively identifying and leveraging deeper latent knowledge, it enhances the expressiveness and learning potential of LLMs.
+Through the coordinated integration of targeted exploration, online filtering, and experience replay, EFRame not only addresses the limitations of GRPO but also establishes a principled and controllable training paradigm.
+Empirical results confirm its efficacy: on the challenging Geometry3K benchmark, EFRame outperforms GRPO by 37.9%, and it achieves consistent improvements across diverse mathematical and multimodal reasoning benchmarks, demonstrating strong robustness and generalizability.
+EFRame provides a systematic solution to the inherent trade-offs among training stability, efficiency, and exploration capacity in RL.
 By enabling the model to effectively identify and leverage deeper latent knowledge, EFRame enhances the expressiveness and learning potential of large language models.
 Through the coordinated integration of targeted exploration, online sample filtering, and experience replay, the framework not only addresses the limitations of GRPO in handling complex reasoning tasks but also establishes a more principled and controllable training paradigm.
-Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 36.5%, and it achieves consistent gains across a wide range of multimodal and mathematical reasoning benchmarks, demonstrating strong generalizability and robustness.
+Empirical evaluations confirm the efficacy of EFRame: on the challenging Geometry3K benchmark, it outperforms GRPO by 37.9%, and it achieves consistent gains across a wide range of mathematical and multimodal reasoning benchmarks, demonstrating strong generalizability and robustness.
 2 Related Work
@@
 As a lightweight variant of PPO, GRPO Shao et al.
-(2024) significantly improves training efficiency, but at the cost of increased instability.
+(2024) significantly improves training efficiency, but at the cost of limited exploration and training instability.
 A series of subsequent studies have introduced optimizations from GRPO.
-These include gradient loss refinements, as demonstrated in DAPO Yu et al.
+DAPO Yu et al.
 (2025) and Dr.GRPO Liu et al.
-(2025); online filtering methods like Meng et al.
-(2025); advantage enhancement techniques, exemplified by VL-Rethinker Wang et al.
-(2025a) and Luffy Yan et al.
-(2025); clip acceleration methods such as CPPO Lin et al.
-(2025); curriculum learning strategies explored in Chen et al.
-(2025b; c); and unsupervised RL such as Zhang et al.
-(2025); Wei et al.
-(2025a); Zhao et al.
-(2025); Agarwal et al.
-(2025).
-While existing methods typically focus on individual components, we systematically review the fundamental limitations of GRPO and propose a comprehensive framework with an integrated solution that addresses three critical aspects: exploration capacity, training stability, and sample efficiency.
-This helps create a more robust and effective GRPO-based training paradigm.
-3 Overview of GRPO
-3.1 Gradient Loss
-GRPO has been employed to enhance the reasoning capabilities of large language models, particularly in mathematical problem-solving.
-Compared to PPO, GRPO improves computational efficiency by eliminating the need for a separate critic model.
-Instead, it estimates the advantage directly using group-normalized rewards.
+(2025) adjust the policy gradient from a more global perspective, but they do not directly address the aforementioned issues.
+Entropy-driven exploration methods, e.g., Cheng et al.
+(2025); Chen et al.
+(2025), introduce entropy regularization into reward or advantage functions to encourage exploration.
+However, this strategy only provides a coarse trend, making entropy difficult to control precisely and stably, while also introducing biases that are hard to estimate.
+Instead, we adopt an exploration method with limited bias—high quality sampling—which enables more effective exploration with controlled deviation.
+Online filtering accelerations, such as Meng et al.
+(2025); Lin et al.
+(2025), further strengthen the efficiency advantage of GRPO, but their contribution to training effectiveness is almost negligible.
+Experience replay approaches, including VL-Rethinker Wang et al.
+(2025), RLEP Zhang et al.
+(2025), and RePO Yan et al.
+(2025), enhance GRPO from the perspective of convergence stability.
+However, experience replay may also aggravate the entropy collapse problem in GRPO, and indiscriminate use can hinder the model’s ability to explore more deeply.
+EFRame addresses these challenges by encouraging exploration through conditional sampling, accelerating training with online filtering, and achieving stable convergence via experience replay, thereby delivering consistent improvements to GRPO in exploration, stability, and convergence.
+Figure 2: The overall workflow of EFRame builds upon GRPO by introducing three key components: additional rollout, online filter, and experience replay.
+3 Methodology
+To resolve the limited exploration and training instability in GRPO, we propose the EFRame, which consists of three main components: additional rollout, online filtering, and experience replay, as illustrated in Figure 2.
+3.1 GRPO and Regular Rollout
+Group Relative Policy Optimization (GRPO) directly estimates the advantage using group-normalized rewards.
 Given a question $q$ and a corresponding verifiable answer $a$, GRPO generates a set of responses $O=\{o_{i}\}_{i=1}^{G}$ using the policy model $\pi_{\theta_{old}}$, then optimizes the policy model by maximizing the following objective:
-$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{(q,a)\sim\mathcal{D},\ O=\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}(\cdot|q)}}\\ &\frac{1}{G}\sum_{i=1}^{G}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t},\ \operatorname{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]-\beta D_{KL}[\pi_{\theta}||\pi_{ref}]\right\},\end{split}$ (1)
-where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i}<t)}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i}<t)}$ and $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G})}$.
-3.2 Limitations of GRPO
-Although GRPO has demonstrated strong performance in various settings, it encounters several practical challenges when applied to complex reasoning tasks.
-In the following discussion, we highlight three main issues: limited exploration capacity, training collapse, and low sample efficiency.
-These limitations reveal the need for more robust and adaptive RL strategies.
-Limited Exploration Capacity.
-GRPO lacks effective mechanisms to promote exploration, leading to persistently low entropy levels in the early stages of training and an increased risk of convergence to local optima.
-Moreover, it does not differentiate between prompts of varying difficulty, which limits its ability to stimulate deeper reasoning capabilities.
-As a consequence, only superficial knowledge is activated, while deeper knowledge remains underutilized.
-((a))
-((b))
-((c))
-Figure 1: Increasing the sampling number to $G=20$ provides a performance improvement over $G=5$, but it incurs a fourfold increase in per-step runtime, leading to very low sample efficiency in practice.
-Low Sample Efficiency.
-GRPO computes sample advantage based on relative advantage.
-When the number of rollouts is small, it suffers from a severe issue of advantage vanishing, leading to a lack of effective training signals.
-However, as the number of samples increases, the relative advantage becomes diluted, making it harder for high-quality samples to exert sufficient influence.
-This results in low sample efficiency: although the computational overhead grows nearly linearly with the number of samples, the corresponding performance improvement is significantly diminished, as shown in Fig.
-1.
-Instability.
-During the later stages of training, the GRPO algorithm may exhibit signs of instability or collapse, characterized by a sharp decline in reward and sudden spikes in both gradient and policy entropy, as illustrated in Fig.
-2.
-In extreme cases, this can escalate into a gradient explosion.
-Such instability presents a serious challenge to reliable long-horizon optimization, particularly in complex reasoning tasks where consistent gradient signals are critical for convergence.
-One contributing factor is the increased variance caused by the disappearance of relative advantages in the later stages of training.
-Supplementing with additional samples with non-zero advantages can help alleviate this issue.
-((a))
-((b))
-((c))
-Figure 2: When training Qwen2.5-VL-7B-Instruct on the Geo3k dataset using the GRPO algorithm, a collapse occurred.
-4 Methodology
-To address the issues in the previous section, we propose the EFRame, which consists of three main components: additional rollout, online filtering and replay buffer, as illustrated in Figure 3.
-Figure 3: The overall workflow of EFRame builds upon the vanilla GRPO framework by introducing three key components: additional rollout, online filtering, and replay buffer.
-4.1 Regular Rollout
-For each batch of prompts, we begin with a regular rollout phase, following the original GRPO training process. This initial rollout serves two key purposes:
+$\begin{split}\mathcal{J}_{GRPO}(\theta)&=\mathbb{E}_{(q,a)\sim\mathcal{D},\ O=\{o_{i}\}_{i=1}^{G_{1}}\sim\pi_{\theta_{old}(\cdot|q)}}\\ &\frac{1}{G_{1}}\sum_{i=1}^{G_{1}}\frac{1}{|o_{i}|}\sum_{t=1}^{|o_{i}|}\left\{\min\left[r_{i,t}(\theta)\hat{A}_{i,t},\ \text{clip}(r_{i,t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{i,t}\right]-\beta D_{KL}[\pi_{\theta}||\pi_{ref}]\right\},\end{split}$ (1)
+where $r_{i,t}(\theta)=\frac{\pi_{\theta}(o_{i,t}|q,o_{i}<t)}{\pi_{\theta_{old}}(o_{i,t}|q,o_{i}<t)}$ and $\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{1}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{1}})}$.
+In EFRame, training begins with a regular rollout phase following this GRPO formulation. This initial rollout serves two key purposes:
   • Obtain reward signals from regular (i.e., solvable) problems.
-  • Identify more challenging prompts that fail to yield positive samples, which will be handled in subsequent stages.
-Valuable samples are temporarily stored in the buffer for gradient updates.
+  • Identify more challenging prompts that fail to yield positive samples, which will be addressed in subsequent stages.
 3.2 Additional Rollout
@@
 Definition 3.1.
-In GRPO-based algorithms with binary rewards (Premise, omitted in the following), difficult prompts are defined as those for which all sampled responses within the initial $G_{1}$ rollouts ($G_{1}$ is a smaller integer, such as 5) fail to obtain a positive reward, where $\bar{\mathcal{D}}\subseteq\mathcal{D}$ and:
-$\bar{\mathcal{D}}=\left\{(q,a)~{}:~{}|\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}~{}:~{}\text{is\_equivalent}(a,o_{i})\right\}_{i=1}^{G_{1}}|=0\right\},$
+In GRPO-based algorithms with binary rewards (Premise, omitted in the following), difficult prompts are defined as those for which all sampled responses within the initial $G_{1}$ rollouts ($G_{1}$ is a small integer, such as 5) fail to obtain a positive reward, where $\bar{\mathcal{D}}\subseteq\mathcal{D}$ and:
+$\bar{\mathcal{D}}=\left\{(q,a)~:~|\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}~:~\text{is\_equivalent}(a,o_{i})\right\}_{i=1}^{G_{1}}|=0\right\}.$
 The additional rollout is designed to activate deeper reasoning capabilities and increase the frequency of emergent “aha moments.”
@@
 While additional rollout enhances the model’s ability to explore complex prompts, it also introduces computational overhead and potential training noise.
-In addition to filtering overlong samples, we design two additional filtering strategies to further enhance the stability and efficiency of the training process.
+In addition to filtering overlong samples introduced in DAPO Yu et al.
+(2025), we design two additional filtering strategies to further enhance the stability and efficiency of the training process.
 Regular Filter: In the first-round rollout, when the prompt is of regular difficulty, we adopt a conservative filtering strategy that retains all regular samples $O_{R}$ to balance efficiency and performance.
@@
 High-quality samples $O_{H}$ are positive samples for difficult prompts $\bar{\mathcal{D}}$:
-$O_{H}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}>0\right\};$
+$O_{H}=\left\{\{o_{i}\}\sim\pi^{t_{a}}_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}>0\right\};$
 low-quality samples are negative samples for difficult prompts $\bar{\mathcal{D}}$:
-$O_{L}=\left\{\{o_{i}\}\sim\pi_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}<0\right\}.$
+$O_{L}=\left\{\{o_{i}\}\sim\pi^{t_{a}}_{\theta_{old}(\cdot|q)}:(q,a)\sim\bar{\mathcal{D}},\hat{A}_{i,t}=\frac{R_{i}-\operatorname{mean}(\{R_{i}\}_{i=1}^{G_{2}})}{\operatorname{std}(\{R_{i}\}_{i=1}^{G_{2}})}<0\right\},$
+where $\pi^{t_{a}}_{\theta_{old}}$ denotes the distribution of $\pi_{\theta_{old}}$ under temperature $t_{a}$. In our setting, $t_{a}=1.2$ by default.
 Theorem 3.4.
@@
 To summarize, the policy gradient under our framework can be expressed as:
-[…]
-$\qquad\qquad\quad{}-\,\beta\,D_{\mathrm{KL}}\bigl{[}\pi_{\theta}\,\|\,\pi_{\mathrm{ref}}\bigr{]}\Bigr{)}\Biggr{]}.$
-To prevent token-length bias, we compute the policy loss using the token-mean aggregation strategy.
-4.4 Replay buffer
+$\mathcal{J}(\theta)$ $=\mathbb{E}\Biggl[\frac{1}{\displaystyle\sum_{o_{i}\in O_{R}\cup O_{H}}\lvert o_{i}\rvert}\sum_{o_{i}\in O_{R}\cup O_{H}}\sum_{t=1}^{\lvert o_{i}\rvert}\Bigl(\min\bigl\{\,r_{i,t}(\theta)\,\hat{A}_{i,t},\;\mathrm{clip}\bigl(r_{i,t}(\theta),1-\epsilon_{\mathrm{low}},1+\epsilon_{\mathrm{high}}\bigr)\,\hat{A}_{i,t}\bigr\}\Biggr].$ (2)
+To prevent token-length bias, we compute the policy loss using a token-mean aggregation strategy.
+Furthermore, following the conclusion of DAPO, and given that catastrophic forgetting is relatively minor in RL Lai et al.
+(2025), we omit KL regularization during training to allow a more thorough optimization of the model.
+3.4 Experience replay
 While GRPO is fundamentally an online RL algorithm, we find that experience replay with high-quality offline samples remains highly effective.
@@
 By incorporating these samples into the replay buffer, their probability of reoccurrence is effectively increased during training, enabling the model to transition from entropy increase to eventual convergence.
-In fact, the replay buffer here serves to reduce policy entropy; detailed discussions can be found in the following discussion section.
-4.5 Discussion
-The above subsections yield several practical insights regarding the interplay between Exploration-Filter-Replay:
-Additional rollout: By prompting LLM to generate multiple responses to challenging queries ($x\in\bar{\mathcal{D}}$) under a high-temperature setting, the resulting positive samples are obtained with extremely low probability $\log\pi_{\theta}(y\mid x)$ and high advantage $\hat{A}(x,y)$.
-According to Theorem 4.5, this high-advantage, low-probability scenario rapidly increases the policy entropy $H$, thereby preventing the collapse of the entropy and maintaining effective exploration.
-At the same time, an excessive presence of low-quality samples can cause a nosedive in entropy, which is difficult to control.
-To address this, we filter out such samples and instead adopt a more controllable strategy to regulate entropy stability.
-Replay Buffer: By storing high-quality positive samples within the replay buffer, the model can repeatedly utilize their gradient information for subsequent updates.
-As the model has already been updated on these (problem, answer) pairs $(x,y)$, the recalculated $\log\pi_{\theta}(y\mid x)$ becomes larger upon rehearing these pairs.
-According to the theorem, this shift leads to a reduction in policy entropy, preventing excessive growth in entropy.
+In fact, the experience replay here serves to reduce policy entropy; detailed discussions can be found in the following subsection.
 4 Experiments
-To validate the effectiveness of our methods, we demonstrate our experiment setup and the results in the following sections.
-Table 1: Comparative results across multiple multimodal reasoning benchmarks indicate that our method achieves excellent performance.
-Models Size MathVision MathVerse MathVista We-Math Average
-Close-Source Models
-GPT-4o Hurst et al. (2024) - 33.95 48.83 59.50 65.00 51.82
-Claude-3.5-Sonnet Anthropic (2024) - 46.48 57.64 68.20 73.05 61.34
-Open-Source Multi-Modal Large Reasoning Models
-QvQ-Preview Bai et al. (2025) 72B 35.56 52.81 69.20 65.29 55.72
-Skywork R1V Peng et al. (2025) 38B 39.31 40.91 67.50 60.06 51.95
-Open-Source 7B Multi-Modal Models
-MM-Eureka Meng et al. (2025) 7B 28.36 49.52 70.40 66.03 53.58
-OpenVLThinker Deng et al. (2025) 7B 25.30 42.79 64.10 63.91 49.03
-ThinkLite-VL Wang et al. (2025b) 7B 27.71 47.41 69.00 67.99 53.03
-VLAA-Thinker Chen et al. (2025a) 7B 26.61 47.49 66.20 66.49 51.70
-Qwen2.5-VL-7B-Instruct with RL
-Qwen2.5-VL Bai et al. (2025) 7B 24.87 43.83 66.30 62.87 49.47
-Qwen2.5-VL+GRPO 7B 29.11 47.51 72.60 67.53 54.19
-Qwen2.5-VL+DAPO 7B 27.92 48.48 72.30 69.08 54.45
-Ours 7B 28.82 (+3.95) 48.81 (+4.98) 73.40 (+7.10) 69.48 (+6.61) 55.13 (+5.66)
+To validate the effectiveness of our methods, we present the experimental setup and results in the following sections.
+Our work is based on the EasyR1 and VeRL frameworks Yaowei et al.
+(2025); Sheng et al.
+(2025), and we compare with the RL baselines GRPO Shao et al.
+(2024) and its improved variant DAPO Yu et al.
+(2025).
+See implementation details in Appendix A.
+Table 1: Comparative results across text-only (LLM) and multimodal (VLM) reasoning benchmarks.
+Text Benchmarks AIME’24 MATH GSM8K Minerva Average
+Qwen2.5-math-7B
+Yang et al. (2024)
+13.3 65.5 65.4 11.0 38.8
+GRPO Shao et al. (2024) 26.7 77.6 87.1 29.0 55.1
+DAPO Yu et al. (2025) 30.0 78.3 87.6 34.2 57.5
+EFRame (ours) 36.7 (+23.4) 80.4 (+19.9) 87.9 (+22.5) 35.7 (+24.7) 60.2 (+21.4)
+Multimodal Benchmarks MathVision MathVerse MathVista We-Math Average
+Qwen2.5-VL-7B-Instruct
+Bai et al. (2025)
+24.87 43.83 66.30 62.87 49.47
+GRPO Shao et al. (2024) 29.11 47.51 72.60 67.53 54.19
+DAPO Yu et al. (2025) 27.92 48.48 72.30 69.08 54.45
+EFRame (ours) 28.82 (+3.95) 48.81 (+4.98) 73.40 (+7.10) 69.48 (+6.61) 55.13 (+5.66)
 4.1 Experimental setup
 Training Datasets.
-Our experiments are conducted on two datasets to evaluate the effectiveness of the algorithm in both in-domain and out-of-domain settings: (1) Geometry3K focused on geometric problems Lu et al.
-(2021a), (2) a 6k out-of-domain dataset randomly sampled from the dataset proposed in  Wei et al.
-(2025b; a), which randomly sampled from established open-source resources, including Geometry3K Lu et al.
+Our experiments are conducted on three datasets to evaluate the effectiveness of the algorithm across different domains and tasks: (1) DAPO-17K, a large-scale mathematics out-of-domain (ood) training dataset designed to evaluate the algorithm’s performance on large language models (LLMs); (2) Multimodal ood dataset of size 6k randomly sampled from the dataset proposed in Wei et al.
+(2025b; a), which is collected from established open-source resources, including Geometry3K Lu et al.
 (2021a), GeoQA Chen et al.
@@
 (2024), and ScienceQA Lu et al.
-(2022a).
+(2022a); (3) Geometry3K, focused on geometric problems Lu et al.
+(2021a), which is used to evaluate the algorithm’s in-domain performance and conduct ablation studies.
 Evaluation Benchmarks.
-Our experiments consist of two parts.
-First, we evaluate the overall performance of the algorithms on four open multimodal math benchmarks: MathVision Wang et al.
+First, we evaluate the algorithm on four text-only reasoning benchmarks—AIME’24 HuggingFaceH4 (2025), MATH Lightman et al.
+(2023), GSM8K Cobbe et al.
+(2021), and Minerva Lewkowycz et al.
+(2022)—which are widely used to assess the problem-solving and reasoning capabilities of large language models (LLMs) across different levels of mathematical difficulty and domain coverage.
+These benchmarks span competition-level problems (AIME’24), advanced mathematics problem sets (MATH), grade-school word problems (GSM8K), and science/mathematics questions from Minerva.
+Second, we evaluate on four open multimodal benchmarks: MathVision Wang et al.
 (2024a), MathVista Lu et al.
@@
 (2024), and We-Math Qiao et al.
-(2024).
-These benchmarks cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty.
-Second, we conduct a detailed analysis of algorithms on the Geometry3K dataset to examine the algorithmic behaviors.
-Baselines.
-We compare our models with a range of advanced multimodal large language models.
-These include GPT-4o Hurst et al.
-(2024) and Claude-3.5-Sonnet Anthropic (2024), which are state-of-the-art closed-source MLLMs.
-QvQ-Preview-72B[2] and Skywork R1V-32B  Peng et al.
-(2025) are multimodal large reasoning models specifically designed for R1-like chain-of-thought capabilities.
-At the 7B scale, we compare with MM-Eureka-7B Meng et al.
-(2025), OpenVLThinker-7B Deng et al.
-(2025), ThinkLite-VL-7B Wang et al.
-(2025b), and VLAA-Thinker-Qwen2.5VL-7B Chen et al.
-(2025a).
-All of these 7B models are based on Qwen2.5-VL and are enhanced for mathematical reasoning via fine-tuning, reinforcement learning, or iterative self-improvement techniques.
-In addition, we compare the performance of GRPO and DAPO on the Qwen2.5-VL-7B-Instruct model to assess the impact of different strategies under consistent model and training settings.
-Implementation details.
-Our work is based on EasyR1 and VeRL frameworks Yaowei et al.
-(2025); Sheng et al.
-(2025).
-We initialize all policy models with Qwen2.5-VL-7B-Instruct and conduct experiments on 8 A800 GPUs (40GB).
-Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout $G_{1}$=5, temperature 1.0, and learning rate $10^{-6}$.
-GRPO and our method adopt $\epsilon_{low}=\epsilon_{high}=0.2$ and a binary reward to avoid relative advantage bias (see detail in Appendix A), while DAPO includes $\epsilon_{high}=0.3$ and an additional length-aware linear penalty (up to 0.5) for responses exceeding 1536 tokens.
-For our method, we perform additional rollouts with $G_{2}=20$, temperature = 1.2; responses with absolute advantage larger than 1 are stored in a replay buffer of size 5120, with experience replay triggered every 5 steps.
-Table 2: The performance comparison on the Geometry3K dataset shows that our method achieves a significant improvement over existing approaches.
-Models Qwen2.5-VL-Instruct Ours GRPO DAPO
-Accuracy 38.44 55.24(+16.80) 50.75 (+12.31) 49.09 (+10.65)
-Time/step(s) - 442($\times$1.35) 327($\times$1.00) 182($\times$0.57)
+(2024), which cover diverse problem types such as geometry, charts, and tables, spanning multiple subjects and levels of visual reasoning difficulty.
+Finally, we conduct a detailed analysis of the algorithms on the Geometry3K dataset to examine in-domain performance and perform ablation studies.
 4.2 Main results
+Models Qwen2.5-VL-7B-Instruct EFRame GRPO DAPO
+Accuracy 38.44 55.41(+16.97) 50.75 (+12.31) 49.09 (+10.65)
 ((a))
@@
 Performance.
-In our experiments, we employ the Exploration-Filter-Replay framework, yielding notable improvements on both out-of-domain and in-domain datasets.
-As shown in Table 1, the proposed method delivers consistently strong results across four multimodal reasoning benchmarks, with an average gain of +5.66 over the base model (Qwen2.5-VL-7B-Instruct), outperforming GRPO by 19.9%.
-On the in-domain Geometry3K dataset, the improvement is even more pronounced: Table 2 reports a +16.80 gain, representing a 36.5% relative advantage over GRPO.
-Exploration.
-As shown in Fig.
-4, our framework maintains the policy entropy around 1, striking a balance between sufficient exploration and exploitation.
-In the later training stages, when the vanilla GRPO tends to suffer from vanishing advantages, our method further increases entropy to encourage continued exploration of informative reward signals.
-Stability.
-Our framework leverages additional rollouts to enrich reward signals, while an online filtering mechanism mitigates training variance by excluding low-quality samples.
-Extensive experiments reveal that the proposed method circumvents collapse issues—occasionally encountered in GRPO and DAPO.
-Furthermore, the steadily increasing response length during training reflects the model’s progressively improved reasoning ability.
-5.3 Entropy Control
-To achieve a well-balanced exploration process, we incorporate entropy control into EFRame by adjusting two key hyperparameters: the additional rollout temperature $t_{a}$ and the replay buffer size $R_{s}$.
-Specifically, increasing the temperature in the additional rollout phase encourages broader exploration, enabling the model to discover low-probability but potentially high-reward trajectories.
-At the same time, we increase the $R_{s}$ to promote convergence and prevent excessive entropy growth.
-Together, these mechanisms allow for a controlled yet sufficiently rich exploration process—ensuring that the model explores meaningfully without destabilizing training due to overly random behavior.
+Our Exploration-Filter-Replay framework achieves substantial improvements in both in-domain and out-of-domain settings for text-only and multimodal reasoning tasks.
+As shown in Table 1, on the text benchmarks, our method attains an average score of 60.2, outperforming the RL baselines GRPO and DAPO by +5.1 and +2.7 average points, respectively.
+The largest absolute gain is observed on AIME’24 (+23.4), demonstrating the framework’s strength on challenging competition-level problems.
+On the multimodal benchmarks, our method reaches an average score of 55.13, outperforming GRPO by +0.94.
+On the in-domain Geometry3K dataset, the gain is even more pronounced: Fig.
+3 reports a +16.97 improvement, corresponding to a 37.9% relative advantage over GRPO.
+Exploration and Entropy Control.
+As shown in Fig.3, our framework maintains batch-level entropy around 1, striking a balance between exploration and exploitation.
+In the later stages of training, when GRPO and DAPO tend to suffer from vanishing advantages, our method keeps the exploration of informative reward signals.
+This is achieved by tuning two hyperparameters: the additional rollout temperature $t_{a}$ and the replay buffer size $R_{s}$.
+A higher $t_{a}$ broadens exploration by sampling low-probability yet high-reward trajectories, while a larger $R_{s}$ promotes convergence and suppresses excessive entropy growth.
+As shown in Fig.4, the two factors exert opposite influences on entropy dynamics, with optimal performance observed when entropy stabilizes at a moderate level.
+Claim 1.
+The combination of high-quality sampling and experience replay enables flexible and efficient entropy control.
+To further improve stability, the framework leverages additional rollouts to enrich reward signals, while an online filtering mechanism reduces training variance by removing low-quality samples, resulting in more robust gradients.
+Extensive experiments confirm that this design effectively avoids the collapse issues observed in GRPO and DAPO.
 ((a))
@@
 ((f))
-Accuracy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
-$R_{s}=5120$ 53.74 52.41 55.24
-$R_{s}=2560$ 52.75 54.41 52.75
-$R_{s}=0$ 52.41 51.08 50.25
-Entropy $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
-$R_{s}=5120$ 0.089 0.305 0.726
-$R_{s}=2560$ 0.091 0.312 2.366
-$R_{s}=0$ 0.092 2.437 4.288
-Figure 5: We report the peak accuracy within the first 200 training steps and the average entropy over the first 150 steps.
-Experiments on the GEO3K dataset using Qwen2.5-VL-7B-Instruct under different settings of additional rollout temperature $t_{a}$ and replay buffer size $R_{s}$ demonstrate that increasing $t_{a}$ enhances exploration, reflected in higher entropy, while enlarging $R_{s}$ promotes convergence, leading to lower entropy.
-Optimal performance is achieved when exploration and convergence are well-balanced.
-We observe in Fig.
-5 that increasing the additional rollout temperature encourages broader exploration, leading to a rise in policy entropy.
-In contrast, enlarging the replay buffer size, which increases the number of high-quality samples reused during training, promotes convergence and results in decreased entropy.
-These two hyperparameters exhibit opposing effects on entropy dynamics, providing a practical mechanism for entropy control within EFRame.
-We find that model performance peaks when the entropy stabilizes around a moderate level.
-Fig.
-5 reports the accuracy under different $t_{a}$ and $R_{s}$, highlighting the impact of entropy control of EFRame on learning efficacy.
-Our experiments demonstrate that entropy control can be achieved with high precision by combining high-temperature positive sampling and experience replay, which are more effective than using entropy-based reward functions.
+Accuracy (Entropy) $t_{a}=1.0$ $t_{a}=1.1$ $t_{a}=1.2$
+$R_{s}=5120$ 53.74 (0.09) 52.41 (0.31) 55.41 (0.73)
+$R_{s}=2560$ 52.75 (0.09) 54.41 (0.31) 52.75 (2.37)
+$R_{s}=0$ 52.41 (0.09) 51.08 (2.44) 50.25 (4.29)
+Figure 4: Peak accuracy (first 200 steps) and average entropy (first 150 steps) on Geometry3K with Qwen2.5-VL-7B-Instruct under varying $t_{a}$ and $R_{s}$.
+Higher $t_{a}$ boosts exploration (entropy↑), larger $R_{s}$ aids convergence (entropy↓), and best performance arises from the balance state of entropy.
 5 Ablation Studies
-To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and replay buffer—are essential for achieving the observed improvements in stability and overall performance.
+To assess the contribution of each component in our framework, we conduct an ablation study to examine whether all three modules—additional rollout, online filtering, and experience replay—are essential for achieving the observed improvements in stability and overall performance.
 5.1 Additional Rollout
-Disabling the additional rollout module exposes the model to a fundamental trade-off inherent in GRPO-based training.
-When the number of samples $G_{1}$ is small, the reward signal is constrained by the base model’s pass@$G_{1}$, limiting the model’s ability to solve challenging problems and suppressing the emergence of “aha moments.”
+[…]
 Claim 2.
-High-quality samples have the potential to activate deeper reasoning abilities within the model.
+High-quality samples can unlock the ability of deeper reasoning in RL.
 As shown in Fig.
-6, the model’s reward on the training set plateaus at a relatively low level, restricting further gains in test accuracy.
-In the early stages of training, the entropy remains low, indicating limited exploration.
-However, once the model hits this performance bottleneck, entropy begins to rise—yet this delayed exploration yields minimal returns, suggesting that the model struggles to escape local optima.
-The additional rollout module is specifically designed to break this situation, enabling targeted deep exploration.
+5, training rewards plateau at low levels, constraining test accuracy.
+Early entropy remains low, reflecting limited exploration; once performance stalls, entropy rises but yields little benefit, indicating difficulty escaping local optima.
+The additional rollout module addresses this by driving targeted deep exploration.
 ((a))
@@
 5.2 Online Filtering
-Removing the online filtering module results in significant instability and inefficiency during training.
-Without filtering, the model treats all samples—regardless of their advantage—as equally important, allowing zero-advantage samples to increase gradient variance and low-quality negative trajectories to distort the learning signal.
-This problem is further exacerbated during the additional rollout phase, where abundant incorrect samples significantly amplify gradient variance, making it difficult for the model to retain meaningful patterns from rare high-quality responses.
-As a result, training becomes slower, more chaotic, and more prone to divergence.
+Removing the online filtering module leads to unstable and inefficient training.
+Without filtering, all samples are treated equally: zero-advantage ones inflate gradient variance, and low-quality negatives distort the learning signal.
+This effect worsens in additional rollouts, where numerous incorrect samples overwhelm rare high-quality responses, making training slower, noisier, and prone to divergence.
 Claim 3.
-Low-quality samples from challenging prompts often mislead optimization, causing the model to converge prematurely to suboptimal solutions and impeding further learning.
+Low-quality samples from challenging prompts mislead optimization, driving premature convergence to suboptimal solutions and preventing further learning.
 As shown in Fig.
-6, during the early stages of training, the inclusion of a large number of negative samples significantly boosts the model’s performance on the training set but yields little improvement on the test set.
-This indicates that most low-quality samples lead the model to converge more quickly to a local optimum, offering little benefit to its generalization ability.
-Moreover, the forward and backward computation of these negative samples incurs substantial overhead, reducing the sample efficiency of the overall training process.
-Online filtering plays a crucial role in maintaining gradient clarity and training stability by selectively preserving informative samples and discarding neutral or low-impact ones.
-6.3 Replay Buffer
-When the replay buffer is disabled, we observe a drastic drop in performance, even falling below that of the vanilla GRPO baseline, as shown in Fig.
-6.
-This counterintuitive result reveals that the abundant high-quality samples generated during the additional rollout phase can actually become detrimental without a proper replay mechanism.
+5, many negative samples initially boost training performance but provide little generalization gain, indicating faster convergence to local optima.
+Their forward and backward computation also adds overhead, lowering sample efficiency.
+Online filtering stabilizes training by preserving informative samples while discarding low-quality ones.
+5.3 Experience replay
+Disabling experience replay causes a sharp performance drop, even below the vanilla GRPO baseline (Fig.
+5).
+This counterintuitive result shows that high-quality samples from additional rollouts can be harmful without a replay mechanism.
 Claim 4.
-While high-quality samples may induce high-entropy exploration, incorporating experience replay enables the model to retain and consolidate useful gradients, thus restoring training stability.
-As high-quality samples typically have low generation probabilities, rewarding them shifts the model’s response distribution toward low probability but correct outputs, disrupting the convergence trajectory and consequently increasing entropy.
-Due to their high advantage values and low sampling probabilities, high-quality samples are more likely to have their gradients clipped during training.
-In the absence of a replay buffer, such valuable samples contribute only short-lived and unstable gradients that are quickly overwritten.
-Instead of reinforcing beneficial behavior, this transient gradient signal increases policy uncertainty, pushing the model into a state of chaotic, high-entropy exploration, ultimately harming both learning stability and performance.
-To address this issue, we leverage experience replay to ensure that high-quality samples continuously provide effective gradients, enabling the model to converge along the new trajectory.
-7 Conclusion and Discussion
-In this work, we investigate the impact of sample quality on RL for large language models, with a particular focus on addressing challenging prompts.
-Our findings show that low-quality incorrect samples—those with low absolute advantage—tend to trap the model in local optima and offer little or even negative contribution to training.
-In contrast, high-quality samples with large advantages play a critical role in guiding the model through a meaningful transition from entropy-driven exploration to entropy-reducing convergence.
-This process is essential for eliciting deeper reasoning capabilities within the model.
-To this end, we propose an Exploration-Filter-Replay framework that emphasizes the discovery, selection, and sustained utilization of high-quality samples.
-Our method introduces additional rollouts to enhance reward density, employs online filtering to suppress noisy samples, and integrates experience replay to amplify the impact of high-quality samples across training iterations.
-This strategy not only helps transform isolated “aha moments” into a sustained “aha state” but also shifts reinforcement learning from a “reinforcement lottery” toward a more structured and reliable process of reinforcement learning.
+Although high-quality samples drive high-entropy exploration, experience replay preserves their gradients, restoring stability and performance.
+High-quality samples usually have low generation probabilities: rewarding them shifts responses toward rare but correct outputs, raising entropy, and disturbing convergence.
+Their large advantage yet low probability also makes their gradients prone to clipping, leaving only fleeting signals that are quickly overwritten.
+Without replay, this transient effect amplifies uncertainty and destabilizes training.
+Experience replay ensures these rare samples repeatedly contribute effective gradients, guiding convergence along a stable trajectory.
+6 Conclusion
+We examined the limitations of GRPO on complex reasoning tasks and identified two major issues: limited exploration and training instability.
+To address these, we proposed EFRame, an Exploration-Filter-Replay framework that enriches exploration with additional rollouts, stabilizes training via online filtering, and improves convergence by replaying rare but informative samples.
+This cycle transforms RL from a “reinforcement lottery” into a more structured and controllable process.
+Ablation studies show that high-advantage positives are crucial for deeper reasoning, while low-advantage negatives drive premature convergence, highlighting the value of selective filtering.
+EFRame further enables fine-grained entropy control by jointly tuning rollout temperature and replay ratios, avoiding biases of entropy regularization and mitigating collapse.
+Future work will focus on adaptive balancing of exploration and replay, more principled filtering, and integration with verifiable rewards and workflow-based training, paving a path toward more reliable and scalable reasoning in LLMs.
+Declaration of AI
+AI is only used for translation and language polishing in this paper.
 Appendix
-Appendix A Relative Advantage Bias of GRPO
-In GRPO, computing relative advantages within a group can introduce bias, especially when the reward becomes multi-dimensional.
-For example, as shown in Fig.
-7, if the reward comprises both accuracy and auxiliary signals such as length, format, or entropy, a significant issue arises when the accuracy signal vanishes within a group—i.e.
-when all responses are either correct or incorrect.
-In such cases, auxiliary rewards (e.g., length) dominate the advantage computation.
-This may result in incorrect responses being rewarded simply for being shorter, or correct responses being penalized for being longer.
-Such misleading gradients introduce irreversible noise into training and lead to reward hacking.
-Therefore, in the training involved in this paper, we use binary rewards whenever possible.
-Figure 7: During training with GRPO, when the primary reward signal (accuracy) vanishes, other auxiliary rewards become dominant, which introduces bias.
+Appendix A Implementation details
+We initialize all policy models and conduct experiments on 8 A800 GPUs (40GB).
+Unless otherwise specified, we follow the default EasyR1 settings: maximum response length of 2048, global batch size 128, rollout batch size 512, rollout $G_{1}$=5, temperature 1.0, and learning rate $10^{-6}$.
+GRPO and our method adopt $\epsilon_{low}=\epsilon_{high}=0.2$ and a binary reward, while DAPO includes $\epsilon_{high}=0.3$ and an additional length-aware linear penalty (up to 0.5) for responses exceeding 1536 tokens.
+For our method, we perform additional rollouts with $G_{2}=20$, temperature = 1.2; responses with absolute advantage larger than 1 are stored in a replay buffer of size 5120, with experience replay triggered every 5 steps.
+Appendix B Limitation
+Although EFRame demonstrates substantial improvements across diverse benchmarks, it imposes a non-trivial requirement on the difficulty level of the training dataset.
+Specifically, EFRame benefits most when the dataset maintains a moderate difficulty level.
+If the dataset is overly challenging, both the regular rollout and the additional rollout stages suffer from excessively high failure rates, resulting in sparse reward signals.
+Conversely, if the dataset is too simple, the proportion of difficult prompts is low, rendering the additional rollout phase largely ineffective due to the scarcity of hard problems.
+In this work, the selected datasets—DAPO-17K, the multimodal open-domain dataset, and Geometry3K—satisfy this moderate-difficulty criterion, enabling the additional rollout and filtering mechanisms to operate effectively.
+However, when training on the MATH-12K dataset, EFRame failed to deliver noticeable gains, with performance being nearly indistinguishable from the GRPO and DAPO baselines.
+This degradation is attributed to the dataset’s excessive simplicity, which results in an insufficient number of high-quality samples discovered during additional rollout exploration.
+This observation underscores that the success of EFRame is contingent on the availability of a sufficient proportion of moderately difficult samples to balance exploration efficacy and training stability.